One mole of monatomic gas is heated at constant pressure of 1 ATM from 0 kelvin to 100 Kelvin if the gas constant is equal 8.32 j/mol/k the change in internal energy of the gas is approximately

Grade:12th pass

1 Answers

Yasaswini Reddy
5 years ago
∆U=nCv dTHere,since n=1,∆U=Cv dTCv=(R)/(gamma-1)Gamma value for a monoatomic gas is 3/2So,Cv=3/2RdT∆U = 1×3/2×8.32×100 =1248 ~=1.25×10^3Hope it helps!Best regards,Yasaswini Reddy 😉